# A Finite-Sample Strong Converse for Binary Hypothesis Testing via (Reverse) Rényi Divergence

### v1

@@ I Introduction
Binary hypothesis testing constitutes a fundamental problem in statistical inference. In the standard setting, an observer collects $n$ independent and identically distributed (i.i.d.) samples from an unknown probability measure $P$. The objective is to decide between two hypotheses: the null hypothesis ($H_{0}$), which posits that $P=P_{0}$, and the alternative hypothesis ($H_{1}$), which posits that $P=P_{1}$.
Binary hypothesis testing finds widespread application in various domains, ranging from signal processing and communications [16, 31] to cybersecurity [8, 25] and machine learning [6, 15], to name a few.
In binary hypothesis testing, depending on the decision, two types of error may occur: the Type I error (or false alarm) which consists of incorrectly rejecting the null hypothesis $H_{0}$ when it is true; and the Type II error (or missed detection), which corresponds to failing to reject $H_{0}$ when the alternative hypothesis $H_{1}$ holds. Ideally, one aims to minimize both probabilities of error; however, a fundamental trade-off exists between these two quantities, preventing the arbitrary reduction of one error probability without affecting the other, and thus making their independent simultaneous minimization unfeasible in many cases.
To address this, a common approach that corresponds to the classical Neyman-Pearson formulation is to minimize the Type II error probability subject to a constraint on the Type I error probability. In this setting, the Neyman-Pearson lemma [23] identifies the Log-Likelihood Ratio Test as the optimal decision rule, achieving the best trade-off possible among the two errors for any sample size.
In the asymptotic regime, as the number of samples $n$ tends to infinity, the fundamental limit of the trade-off between the two probabilities of error is characterized by the Chernoff-Stein lemma [10, 9, 11], which establishes that the optimal error exponent for the Type II error is the Kullback-Leibler (KL) divergence, $D(P_{0}\|P_{1})$, when the Type I error is bounded by a constant. This result has been successively refined and extended to the finite blocklength regime by Espinosa et al. [12], who showed that this optimal rate holds even when the Type I error is not bounded by a constant but decays sub-exponentially with $n$. Conversely, when the Type I error is constrained to decay exponentially with a rate $c$, Blahut [7] and, Han and Kobayashi [14] identified $D(P_{1}\|P_{0})$ as the maximal achievable rate under which the Type II error can asymptotically vanish.
Although these results provide crucial information about the problem structure, the majority only provide asymptotic gaurantees. For this reason, recent research has increasingly shifted focus toward the finite samples regime, employing various information-theoretic measures to derive converse and achievability bounds [27, 3, 26, 20].
@@
A similar characterization of the error can also be established for the $E_{\gamma}$ divergence, which represents a natural generalization of the $\mathrm{TV}$ distance [2, 22].
However, both $\mathrm{TV}$ and $E_{\gamma}$ suffer from a critical limitation: they do not tensorize. As a result, the corresponding bounds for product distributions become computationally intractable as the sample size $n$ increases, as shown by Bhattacharyya et al. [5, 4].
@@
Along a different line, Bar-Yossef [3] established bounds on the Type II error probability in terms of the Squared Hellinger distance between $P_{0}$ and $P_{1}$ and the sample size $n$. Similarly, Pensia et al. [26] focused on the Bayesian setting, where the hypotheses are endowed with prior probabilities and the objective is to minimize the average error probability, and provided a rigorous characterization of the sample complexity—the minimum number of observations required to meet specific error constraints—in terms of Hellinger distances as well. However, these Hellinger-based approaches are inherently symmetric and therefore do not capture the asymmetric nature of binary hypothesis testing. Moreover, as with KL-based approaches, the resulting bounds yield only weak converse guarantees.
These limitations motivate the need for alternative information measures capable of characterising strong converse behaviour in finite sample regimes, and capturing the phase transition between the achievable and impossible regimes. In pursuing this goal, our analysis led us to focus on the Rényi divergence [30], which has previously been shown to be closely related to hypothesis testing and strong converse phenomena [21, 28].
@@ I-A Contributions
Furthermore, a key aspect of our approach relies on the use of the “reverse” Rényi divergence, $D_{\lambda}(P_{1}\|P_{0})$, rather than the more common “forward” direction from $P_{0}$ to $P_{1}$. This also fundamentally distinguishes our approach from those based on symmetric measures, such as the $\mathrm{TV}$ distance and the Hellinger distance.
@@ II Preliminaries and notation
The Neyman-Pearson lemma [23] states that the optimal solution to (2) is achieved by a Log-Likelihood Ratio Test (LLRT). We recall that a LLRT with threshold $\tau\in\mathbb{R}$ is a decision rule $\phi:\mathcal{X}^{n}\to\{0,1\}$ such that
$\phi(\mathbf{x)}=\begin{cases}1&\text{if }\log\frac{P_{1}^{n}(\mathbf{x})}{P_{0}^{n}(\mathbf{x})}\geq\tau,\\ 0&\text{otherwise.}\end{cases}$ (3)
@@ III-A Converse Bounds
In contrast, the second bound, stemming from the “forward” divergence, $D_{\lambda}(P_{0}^{n}\|P_{1}^{n})$, ensures a non-trivial bound even in settings where the first one may become loose ( i.e., negative). For instance, when $\varepsilon$ is fixed and the sample size $n\to\infty$, the inequality $\log(1/\varepsilon)<nD{\lambda}(P_{1}\|P_{0})$ holds for any $\lambda>1$. As a result, in the first bound the subtracted term grows with nnm rendering the bound vacuous for large $n$. In contrast, for a fixed $\varepsilon$, the second bound yields a lower bound on $\beta_{n}(\varepsilon)$ of the form $e^{-nD_{\lambda}(P_{0}\|P_{1})}$.
@@
Corollary 1 provides a necessary condition for the minimum sample size. In other words, given a fixed constraint $\varepsilon$ on the Type I error, if we require the Type II error to be no greater than a chosen threshold $\delta$, the number of samples $n$ must be at least as large as the value defined in (1).
Furthermore, for small values of $\varepsilon$ and $\delta$, the necessary condition on the sample size simplifies to
$\Omega\left(\max\left\{\frac{\log(1/\varepsilon)}{D_{\lambda}(P_{1}\|P_{0})},\frac{\log(1/\delta)}{D_{\lambda}(P_{0}\|P_{1})}\right\}\right).$
@@ III-B Achievability Bound
Let $P_{1}^{n}$ be absolutely continuous with respect to $P_{0}^{n}$. Let $\phi:\mathcal{X}^{n}\to\{0,1\}$ be a LLRT with threshold $\tau\in\mathbb{R}$, and let $\alpha_{\phi}$ and $\beta_{\phi}$ denote the Type I error and Type II error probabilities of $\phi$, respectively. Then, one has
@@ V Comparisons
Beyond characterizing the phase transition between feasible and infeasible regimes, and providing a strong converse result, our bound also yields significant improvements in the finite-sample regime. In this section, we show that our converse bound in Eq. (1) strictly improves upon existing results, yielding tighter lower bounds on the Type II error probability across a variety of settings. For this purpose, we present several numerical experiments focusing on two classical testing scenarios.
@@
For both scenarios, we benchmark our converse bound, under various Type I error constraints, against several well-known lower bounds. Three are based on the KL divergence, which includes Fano’s inequality [13], a bound derived from the Berry-Esseen theorem [27, Lemma 58], and one from the smoothing-out method presented in [18, Thm 2.4]. The other lower bound we compare against is based on the Hellinger distance [3, Thm 4.7]. We exclude the bound obtained through the blowing-up lemma since it is strictly dominated by the smoothing-out bound in the regimes considered. In addition, for each setting, we compute the exact Type II error probability $\beta_{n}(\varepsilon)$ via the Neyman-Pearson Lemma. Figures 1 and 2 illustrate these comparisons under two regimes for the constraint on the Type II error (constant and exponential decay); additional scenarios are considered in Appendix E.
@@
Figure 1: Bernoulli testing ($P_{0}=\text{Bern}(1/2)$ vs $P_{1}=\text{Bern}(1/2+\delta)$, with $\delta=0.01$) under two Type I error ($\varepsilon$) regimes: constant (top) and exponential (bottom). The $x$-axis denotes the sample size $n$, and the $y$-axis the Type II error.
@@
Figure 2: Gaussian testing ($P_{0}=\mathcal{N}(2,1)$ vs $P_{1}=\mathcal{N}(2+\delta,1)$, with $\delta=0.05$) under two Type I error ($\varepsilon$) regimes: constant (top) and exponential (bottom). The $x$-axis denotes the sample size $n$, and the $y$-axis the Type II error.
@@ Appendix A Proof of Theorem 1
We remark that the lower bound can be established via both the variational representation of the Rényi divergence and the Data Processing Inequality (DPI). In the analysis that follows, we present the proof utilizing the DPI.
@@ Appendix B Proof of Corollary 1
Furthermore, if we demand for a test that achieves a Type II error probability below a target threshold $\delta$ (i.e., $\beta_{n}(\varepsilon)\leq\delta$), we can invert the bound in (15) for a fixed $\lambda$. This yields the following necessary condition on the sample size:
@@ Appendix D Proof of Theorem 3
Let $\delta>0$ be the negative gap such that $c=D_{\lambda}(P_{1}\|P_{0})-\delta$ for some $\lambda\in(0,1)$. We need to show that there exists a test that satisfies the constraint on the Type I error while ensuring that the Type II error vanishes as $n$ increases. For such a purpose, consider a LLRT $\phi$ with a threshold $\tau$ defined as follows
@@
First of all, we have to verify that $\phi$ satisfies the Type I error constraint, i.e., its Type I error $\alpha_{n}\leq e^{-nc}$. Markov’s inequality yields
@@
Appendix E Additional Comparisons

### v2

@@ I Introduction
Binary hypothesis testing constitutes a fundamental problem in statistical inference. In the standard setting, an observer collects $n$ independent and identically distributed (i.i.d.) samples from an unknown probability measure $P$. The objective is to decide between two hypotheses: the null hypothesis ($H_{0}$), which posits that $P=P_{0}$, and the alternative hypothesis ($H_{1}$), which posits that $P=P_{1}$. Binary hypothesis testing finds widespread application in various domains, ranging from signal processing and communications [16, 32] to cybersecurity [8, 26] and machine learning [6, 15], to name a few.
In binary hypothesis testing, depending on the decision, two types of error may occur: the Type I error (or false alarm) which consists of incorrectly rejecting the null hypothesis $H_{0}$ when it is true; and the Type II error (or missed detection), which corresponds to failing to reject $H_{0}$ when the alternative hypothesis $H_{1}$ holds. Ideally, one aims to minimize both probabilities of error; however, a fundamental trade-off exists between these two quantities, preventing the arbitrary reduction of one error probability without affecting the other, and thus making their independent simultaneous minimization unfeasible in many cases. To address this, a common approach that corresponds to the classical Neyman-Pearson formulation is to minimize the Type II error probability subject to a constraint on the Type I error probability. In this setting, the Neyman-Pearson lemma [24] identifies the Log-Likelihood Ratio Test as the optimal decision rule, achieving the best trade-off possible among the two errors for any sample size.
In the asymptotic regime, as the number of samples $n$ tends to infinity, the fundamental limit of the trade-off between the two probabilities of error is characterized by the Chernoff-Stein lemma [10, 9, 11], which establishes that the optimal error exponent for the Type II error is the Kullback-Leibler (KL) divergence, $D(P_{0}\|P_{1})$, when the Type I error is bounded by a constant. This result has been successively refined and extended to the finite blocklength regime by Espinosa et al. [12], who showed that this optimal rate holds even when the Type I error is not bounded by a constant but decays sub-exponentially with $n$. Conversely, when the Type I error is constrained to decay exponentially with a rate $c$, Blahut [7] as well as Han and Kobayashi [14] identified $D(P_{1}\|P_{0})$ as the maximal achievable rate under which the Type II error can asymptotically vanish.
Although these results provide crucial information about the problem structure, the majority only provide asymptotic guarantees. For this reason, recent research has increasingly shifted focus toward the finite samples regime, employing various information-theoretic measures to derive converse and achievability bounds [28, 3, 27, 21].
@@
A similar characterization of the error can also be established for the $E_{\gamma}$ divergence, which represents a natural generalization of the $\mathrm{TV}$ distance [2, 23]. However, both $\mathrm{TV}$ and $E_{\gamma}$ suffer from a critical limitation: they do not tensorize. As a result, the corresponding bounds for product distributions become computationally intractable as the sample size $n$ increases, as shown by Bhattacharyya et al. [5, 4].
@@
Along a different line, Bar-Yossef [3] established bounds on the Type II error probability in terms of the Squared Hellinger distance between $P_{0}$ and $P_{1}$ and the sample size $n$. Moreover, Pensia et al. [27, 17] showed that the asymmetric and Bayesian settings, where the hypotheses are endowed with prior probabilities and the objective is to minimize the average error probability, can be treated equivalently. Thus, focusing on the Bayesian setting, they provided a rigorous characterization of the sample complexity—the minimum number of observations required to meet specific error constraints—in terms of Hellinger/Jensen-Shannon divergences.
We focus on characterizing strong converse behaviour in finite sample regimes, and capturing the phase transition between the achievable and impossible regimes. In pursuing this goal, our analysis led us to the Rényi divergence [31], which has previously been shown to be closely related to hypothesis testing and strong converse phenomena [22, 29].
@@ I-A Contributions
Furthermore, a key aspect of our approach relies on the use of the “reverse” Rényi divergence, $D_{\lambda}(P_{1}\|P_{0})$, rather than the more common “forward” direction from $P_{0}$ to $P_{1}$. This also fundamentally distinguishes our approach from those based on symmetric measures, such as the $\mathrm{TV}$ distance and the Squared Hellinger distance.
@@ II Preliminaries and notation
The Neyman-Pearson lemma [24] states that the optimal solution to (2) is achieved by a Log-Likelihood Ratio Test (LLRT). We recall that a deterministic LLRT with threshold $\tau\in\mathbb{R}$ is a decision rule $\phi:\mathcal{X}^{n}\to\{0,1\}$ such that
$\phi(\mathbf{x)}=\begin{cases}1&\quad\text{if }\log\frac{P_{1}^{n}(\mathbf{x})}{P_{0}^{n}(\mathbf{x})}\geq\tau,\\ 0&\quad\text{otherwise.}\end{cases}$ (3)
@@ III-A Converse Bounds
In contrast, the second bound, stemming from the “forward” divergence, $D_{\lambda}(P_{0}^{n}\|P_{1}^{n})$, ensures a non-trivial bound even in settings where the first one may become loose ( i.e., negative). For instance, when $\varepsilon$ is fixed and the sample size $n\to\infty$, the inequality $\log(1/\varepsilon)<nD{\lambda}(P_{1}\|P_{0})$ holds for any $\lambda>1$. As a result, in the first bound the subtracted term grows with $n$, rendering the bound vacuous for large $n$. In contrast, for a fixed $\varepsilon$, the second bound yields a lower bound on $\beta_{n}(\varepsilon)$ of the form $e^{-nD_{\lambda}(P_{0}\|P_{1})}$.
@@
Remark 1
A lower bound characterizing the sample complexity in the Bayesian setting appears in [17, Thm. 2]. By using the equivalence between the Bayesian and asymmetric settings noted in [27, Claim 4.6], it can be rewritten in our notation as
$n\geq\frac{1}{2}\frac{\lambda_{*}}{1-\lambda_{*}}\frac{\log(1/2\varepsilon)}{D_{\lambda_{*}}(P_{0}\|P_{1})},$ (7)
where $\lambda_{*}=\frac{\log(1/2\delta)}{\log(1/2\delta)+\log(1/2\varepsilon)}\in[0.5,1)$. The range of $\lambda$ in Eq.  (7) is different from the one in Eq. (1). A direct comparison is thus not straightforward: in some regimes, Eq. (7) yields a tighter bound than Eq. (1), while in others the reverse holds.
@@ III-B Achievability Bound
Let $P_{1}^{n}$ be absolutely continuous with respect to $P_{0}^{n}$. Let $\phi:\mathcal{X}^{n}\to\{0,1\}$ be an LLRT with threshold $\tau\in\mathbb{R}$, and let $\alpha_{\phi}$ and $\beta_{\phi}$ denote the Type I error and Type II error probabilities of $\phi$, respectively. Then, one has
@@ V Comparisons
Beyond characterizing the phase transition between feasible and infeasible regimes, and providing a strong converse result, our bound also yields significant improvements in the finite-sample regime. In this section, we show that our converse bound in Eq. (1) strictly improves upon existing results, yielding tighter lower bounds on the Type II error probability across a variety of settings. For this purpose, we present several numerical experiments focusing on two classical testing scenarios. Note that our comparisons focus on the Chernoff-Stein regime, which is different from settings where both error types are required to be bounded (e.g., [17]).
@@
For both scenarios, we benchmark our converse bound, under various Type I error constraints, against several well-known lower bounds. Three are based on the KL divergence, which includes Fano’s inequality [13], a bound derived from the Berry-Esseen theorem [28, Lemma 58], and one from the smoothing-out method presented in [19, Thm 2.4]. The other lower bound we compare against is based on the Hellinger distance [3, Thm 4.7]. We exclude the bound obtained through the blowing-up lemma since it is strictly dominated by the smoothing-out bound in the regimes considered. In addition, for each setting, we compute the exact Type II error probability $\beta_{n}(\varepsilon)$ via the Neyman-Pearson Lemma. Figures 1 and 2 illustrate these comparisons under three regimes for the constraint on the Type II error (constant, linear and exponential decay). For the Gaussian example, explicit computations are provided in Appendix E to compare the behavior of the various bounds. In particular, the real Type II error behave as
$\beta_{n}(e^{-nc})\sim 1-\exp\left(-n\left(\sqrt{c}-\sqrt{D(P_{1}\|P_{0})}\right)^{2}\right),$ (12)
with $c>D(P_{1}\|P_{0})$, while Eq. (10) with $\lambda=\sqrt{c/D(P_{1}\|P_{0})}>1$ reduces to
$\beta_{n}(e^{-nc})\geq 1-\exp\left(-n\left(\sqrt{c}-\sqrt{D(P_{1}\|P_{0})}\right)^{2}\right),$ (13)
matching the exact behaviour from Eq. (12). By contrast, the converse bounds existing in the literature that leverage the Kullback-Leibler divergence or Hellinger distance fail to capture the convergence to $1$ of $\beta_{n}$. Additional scenarios of comparison are considered in Appendix F.
@@
Figure 1: Bernoulli testing ($P_{0}=\text{Bern}(1/2)$ vs $P_{1}=\text{Bern}(1/2+\delta)$, with $\delta=0.01$) under three Type I error ($\varepsilon$) regimes: constant (top), linear (middle) and exponential (bottom). The $x$-axis denotes the sample size $n$, and the $y$-axis the Type II error.
@@
Figure 2: Gaussian testing ($P_{0}=\mathcal{N}(2,1)$ vs $P_{1}=\mathcal{N}(2+\delta,1)$, with $\delta=0.05$) under three Type I error ($\varepsilon$) regimes: constant (top), linear (middle) and exponential (bottom). The $x$-axis denotes the sample size $n$, and the $y$-axis the Type II error.
@@ Appendix A Proof of Theorem 1
We remark that the lower bound can be established via both the variational representation of the Rényi divergence and the Data-Processing Inequality (DPI). In the analysis that follows, we present the proof utilizing the DPI.
@@ Appendix B Proof of Corollary 1
Furthermore, if we require a test that achieves a Type II error probability below a target threshold $\delta$ (i.e., $\beta_{n}(\varepsilon)\leq\delta$), we can invert the bound in (18) for a fixed $\lambda$. This yields the following necessary condition on the sample size:
@@ Appendix D Proof of Theorem 3
Let $\delta>0$ be the negative gap such that $c=D_{\lambda}(P_{1}\|P_{0})-\delta$ for some $\lambda\in(0,1)$. We need to show that there exists a test that satisfies the constraint on the Type I error while ensuring that the Type II error vanishes as $n$ increases. For such a purpose, consider an LLRT $\phi$ with a threshold $\tau$ defined as follows
@@
First, we verify that $\phi$ satisfies the Type I error constraint, i.e., its Type I error $\alpha_{n}\leq e^{-nc}$. Markov’s inequality yields
@@
Appendix E Analytical Derivations for testing between Gaussians
In this section, we analyze the behavior of the Type II error probability for a binary hypothesis testing problem between two Gaussian distributions, $P_{0}\sim\mathcal{N}(\mu,1)$ and $P_{1}\sim\mathcal{N}(\mu+\delta,1)$. We specifically consider the regime where the Type I error probability is constrained to decay exponentially fast, i.e., $\alpha_{n}\leq e^{-nc}$, with $c>D(P_{1}\|P_{0})$.
E-A Exact Asymptotic Analysis
By the Neyman-Pearson lemma, the optimal test is a Log-Likelihood Ratio Test (LLRT). The log-likelihood ratio is given by:
$\log\frac{dP_{1}^{n}}{dP_{0}^{n}}(X^{n})$ $=\sum_{i=1}^{n}\left(-\frac{(X_{i}-\mu-\delta)^{2}}{2}+\frac{(X_{i}-\mu)^{2}}{2}\right)$
$=\delta\sum_{i=1}^{n}(X_{i}-\mu)-\frac{n\delta^{2}}{2}.$
The decision rule is equivalent to comparing the sample mean $\bar{X}_{n}$ to a threshold $\gamma_{n}$. Specifically, we reject $H_{0}$ if $\bar{X}_{n}\geq\gamma_{n}$.
The threshold $\gamma_{n}$ is determined by the Type I error constraint. Under $H_{0}$, $\bar{X}_{n}\sim\mathcal{N}(\mu,1/n)$. Thus,
$\alpha_{n}=Q\left(\sqrt{n}(\gamma_{n}-\mu)\right)\leq e^{-nc},$
where $Q(x)=\int_{x}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2}dt$. Assuming that the equality holds, and utilizing the asymptotic approximation $\log Q(x)\approx-x^{2}/2$ for large $x$, we have:
$\frac{n(\gamma_{n}-\mu)^{2}}{2}\approx-nc\implies\gamma_{n}\approx\mu+\sqrt{2c}.$ (25)
The Type II error probability $\beta_{n}$ corresponds to the probability that $\bar{X}_{n}<\gamma_{n}$ under $H_{1}$, where $\bar{X}_{n}\sim\mathcal{N}(\mu+\delta,1/n)$.
$\beta_{n}$ $=P_{1}(\bar{X}_{n}<\gamma_{n})$
$=\Phi\left(\sqrt{n}(\gamma_{n}-(\mu+\delta))\right)$
$=\Phi\left(\sqrt{n}(\gamma_{n}-\mu-\delta)\right).$ (26)
Substituting the approximation from (25) into (26), and using the fact that the KL divergence between the two distributions is given by $D(P_{1}\|P_{0})=\frac{\delta^{2}}{2}$.
$\beta_{n}\approx\Phi\left(\sqrt{n}(\sqrt{2c}-\delta)\right)=\Phi\left(\sqrt{2n}(\sqrt{c}-\sqrt{D(P_{1}\|P_{0})})\right).$
When $c>D(P_{1}\|P_{0})$, the argument is positive. Thus, we analyze how close $\beta_{n}$ is to 1 by considering $1-\beta_{n}=Q(\cdot)$:
$\beta_{n}\sim 1-\exp\left(-n\left(\sqrt{c}-\sqrt{D(P_{1}\|P_{0})}\right)^{2}\right).$ (27)
E-B Converse via Theorem 1
First, we compute the Rényi divergence of order $\lambda$:
$D_{\lambda}(P_{1}\|P_{0})=\lambda\frac{\delta^{2}}{2}=\lambda D(P_{1}\|P_{0}).$ (28)
From Eq. (9) in Theorem 3, when $\alpha_{n}\leq e^{-nc}$ with $c>D(P_{1}\|P_{0}$), for any $n$, the Type II error is lower bounded by:
$\beta_{n}\geq 1-\inf_{\lambda>1}\exp\left(-n\frac{\lambda-1}{\lambda}(c-D_{\lambda}(P_{1}\|P_{0}))\right).$
Substituting (28), the exponent term becomes:
$E(\lambda)\triangleq\left(1-\frac{1}{\lambda}\right)\big(c-\lambda D(P_{1}\|P_{0})\big).$
To find the tightest bound, we maximize $E(\lambda)$ with respect to $\lambda>1$. The derivative is $\frac{d}{d\lambda}E(\lambda)=-D(P_{1}\|P_{0})+\frac{c}{\lambda^{2}}$ and setting it to zero yields $\lambda^{*}=\sqrt{c/D(P_{1}\|P_{0})}$. Since $c>D(P_{1}\|P_{0})$, $\lambda^{*}>1$, which is valid. Evaluating the exponent at $\lambda^{*}$ gives $E(\lambda^{*})=(\sqrt{c}-\sqrt{D})^{2}$ so that the bound becomes
$\beta_{n}\geq 1-\exp\left(-n\left(\sqrt{c}-\sqrt{D(P_{1}\|P_{0})}\right)^{2}\right).$
We observe that this matches the exact behavior of the Type II error computed in Eq. (27).
E-C Converse via Fano’s Inequality
In this case, the lower bound on the Type II error $\beta_{n}$ is defined as follows:
$\beta_{n}\geq\frac{e^{-nD(P_{0}\|P_{1})-\log 2}}{1-\alpha_{n}}.$
Since $\alpha_{n}\leq e^{-nc}$, we obtain:
$\beta_{n}\geq\frac{e^{-nD(P_{0}\|P_{1})-\log 2}}{1-e^{-nc}}.$ (29)
Since $c>D(P_{1}\|P_{0})$, the denominator in (29) approaches 1 as $n\to\infty$. Consequently, the bound asserts:
$\beta_{n}\gtrsim e^{-nD(P_{0}\|P_{1})}.$
However, from Eq. (27) for $c>D(P_{1}\|P_{0})$, the Type II error probability $\beta_{n}$ converges to 1. Thus, the inequality derived via Fano’s Inequality is valid but provides no information about the rate at which $\beta_{n}$ approaches 1 in this setting.
E-D Converse via Hellinger Distance
Let $H^{2}(P_{0},P_{1})$ denote the squared Hellinger distance between $P_{0}$ and $P_{1}$, defined as follows:
$H^{2}(P_{0},P_{1})$ $=1-\int_{\mathcal{X}}\sqrt{p_{0}(\mathbf{x})p_{1}(\mathbf{x})}\,d\mu(\mathbf{x}).$
From the approach presented in [3, Thm 4.7], one derive the following converse bound on the Type II error:
$\beta_{n}\geq 1-\sqrt{1-(1-H^{2}(P_{0},P_{1}))^{2n}}-\alpha_{n}.$
Recalling that $1-H^{2}(P_{0},P_{1})=e^{-\frac{\delta^{2}}{8}}$ and $D(P_{1}\|P_{0})=\frac{\delta^{2}}{2}$, we have that $1-H^{2}(P_{0},P_{1})=e^{-\frac{D(P_{1}\|P_{0})}{4}}$. Moreover, since $\alpha_{n}\leq e^{-nc}$, we obtain:
$\beta_{n}\geq 1-\sqrt{1-e^{-\frac{nD(P_{1}\|P_{0})}{2}}}-e^{-nc}.$ (30)
For $c>D(P_{1}\|P_{0})$, one can see that both the term $e^{-\frac{nD(P_{1}\|P_{0})}{2}}$ and $e^{-nc}$ in Eq. (30) vanishes as $n\to\infty$. Therefore, as Fano’s Inequality, it provides no information about the rate at which $\beta_{n}$ approaches 1.
E-E Converse via Berry-Esseen Theorem
From  [28, Lemma 58], one has the following lower bound on the Type II error:
$\log(\beta_{n})\geq-nD(P_{0}\|P_{1})-\sqrt{nV_{n}}Q^{-1}\left(1-\alpha_{n}-\frac{B_{n}+\Delta}{\sqrt{n}}\right)+\log\Delta-\frac{1}{2}\log n,$
where $V_{n}=\text{Var}_{P_{0}}\left[\log\frac{p_{0}(\mathbf{X})}{p_{1}(\mathbf{X})}\right]$, $T_{n}=\mathbb{E}_{P_{0}}\left[\left|\log\frac{p_{0}(\mathbf{X})}{p_{1}(\mathbf{X})}-D(P_{0}\|P_{1})\right|^{3}\right]$, $B_{n}=6\frac{T_{n}}{V_{n}^{3/2}}$ and $\Delta>0$.
For the Gaussian case $P_{0}\sim\mathcal{N}(\mu,1)$ and $P_{1}\sim\mathcal{N}(\mu+\delta,1)$, we have that $V_{n}=\delta^{2}$, $T_{n}=|\delta|^{3}\sqrt{8/\pi}$ and $B_{n}=6\sqrt{8/\pi}$. Thus, by fixing $\Delta>0$ we obtain:
$\log(\beta_{n})\geq-nD(P_{0}\|P_{1})-\sqrt{n\delta^{2}}Q^{-1}\left(1-\alpha_{n}-\frac{6\sqrt{8/\pi}+\Delta}{\sqrt{n}}\right)+\log\Delta-\frac{1}{2}\log n.$
As $n\to\infty$, since $\alpha_{n}\leq e^{-nc}$, the argument of $Q^{-1}(\cdot)$ approaches 1. Using the approximation $Q^{-1}(1-\epsilon)\sim-\sqrt{2\log(1/\epsilon)}$, with $\epsilon=\alpha_{n}+\frac{6\sqrt{8/\pi}+\Delta}{\sqrt{n}}$, we get
$\log(\beta_{n})\geq-nD(P_{0}\|P_{1})+\sqrt{n\delta^{2}}\sqrt{\log n-2\log(\sqrt{n}\alpha_{n}+6\sqrt{8/\pi}+\Delta)}+\log\Delta-\frac{1}{2}\log n.$ (31)
As $n\to\infty$, since $\alpha_{n}\leq e^{-nc}$ for $c>D(P_{1}\|P_{0})$, the second term in (31) scales as $\sqrt{\delta^{2}n\log n}$. Thus, recalling that $D(P_{1}\|P_{0})=\delta^{2}/2$, it follows that the term $nD(P_{0}\|P_{1})$ dominates the expression (31). This implies that for $c>D(P_{1}\|P_{0})$, the lower bound decays to 0. In contrast, the exact Type II error actually converges to 1 as shown in (27).
E-F Converse via Smoothing Out
Applying the smoothing-out method presented in [19, Thm 2.4] to this Gaussian setting, we have that for any $t>0$ the following bound holds:
$\log(\beta_{n})\geq-nD(P_{0}\|P_{1})+\frac{1}{1-e^{-2t}}\log(1-\alpha_{n})-nt-\frac{\delta^{2}}{2}\left(e^{t}-1\right)^{2}-n\big(\cosh(2t)-1\big),$
where $\cosh(t)=\frac{e^{t}+e^{-t}}{2}$.
To find the tightest bound, we have to observe how the parameter $t$ behaves. If $t$ is fixed, as $n\to\infty$, since $\alpha_{n}\leq e^{-nc}$, the terms nt-nt and $n(\cosh(2t)-1)$ dominate the lower bound, forcing it to go to -$\infty$ (i.e., $\beta_{n}\geq 0$). Thus, the parameter $t$ must vanish as $n$ increases. Using the Taylor expansions $e^{-2t}\approx 1-2t$, $\log(1-x)\approx-x$ for small $x$, and $\cosh(2t)-1\approx 2t^{2}$, the bound simplifies to:
$\log(\beta_{n})\gtrsim-nD(P_{0}\|P_{1})-\frac{\alpha_{n}}{2t}-nt-2nt^{2}.$
Optimizing for $t$ yields $t\approx\sqrt{\alpha_{n}/2n}$. However, because $\alpha_{n}\leq e^{-nc}$, this choice makes the term $nD(P_{0}\|P_{1})$ the dominant one, as $n\to\infty$. Consequently, in the regime $c>D(P_{1}\|P_{0})$, the lower bound fails to capture the fact that the exact Type II error $\beta_{n}$ converges to 1.
Appendix F Additional Comparisons